# The optimal hypercontractive constants for $\mathbb{Z}_{3}$

### v1

@@ 4.1. Existence of a nontrivial extremizer
It is clear that $G(r_{p,q},0)=0$. We shall now show that there exists a nonzero extremizer, which corresponds to a root of $G(r_{p,q},\rho)=0$ in the open interval $(0,1)$.
@@
By the definition of $r_{p,q}$, we have $G(r_{p,q},\rho)\geq 0$ for all $\rho\in[0,1]$. In particular, we have $G(r_{p,q},1)\geq 0$. We now prove that $G(r_{p,q},1)\neq 0$. Suppose by contradiction that $G(r_{p,q},1)=0$. Then the left derivative of $G(r_{p,q},\cdot)$ at $\rho=1$ is non-positive. Therefore,
@@
By Lemma 3.1, we may find and fix a $r_{0}\in(r_{p,q},\sqrt{(p-1)/(q-1)})$. By $r_{0}<\sqrt{(p-1)/(q-1)}$ and the expansion (3.3) of $G(r,\rho)$ at $\rho=0$, there exists $\rho_{1}=\rho_{1}(r_{0})\in(0,1)$ such that $G(r_{0},\rho)\geq 0$ for all $\rho\in[0,\rho_{1}]$. Observe that $\|T_{r}f_{\rho}\|$ is increasing in $r\in[0,1]$, which implies that $G(r,\rho)$ is decreasing with respect to $r$. Hence, for any $r\in[0,r_{0}]$, we have
$G(r,\rho)\geq G(r_{0},\rho)\geq 0,\ \forall\ \rho\in[0,\rho_{1}].$
Then by Proposition 3.2 and the definition of $r_{p,q}$, for any $r\in(r_{p,q},r_{0}]$, we have
$\min\{G(r,\rho)\mid\rho\in[\rho_{1},1]\}<0.$
By the definition of $r_{p,q}$ and the continuity of $G(r,\rho)$, letting $r\to r_{p,q}^{+}$ gives
$0\leq\min\{G(r_{p,q},\rho)\mid\rho\in[\rho_{1},1]\}\leq 0.$
That is, there exists $\rho_{0}\in[\rho_{1},1]$ such that $G(r_{p,q},\rho_{0})=0$. Finally, Lemma 4.2 implies that $\rho_{0}\in[\rho_{1},1)\subset(0,1)$, which finishes the proof. ∎
@@ 4.2. Derivation of the extremal system
To complete the proof of Theorem 1.1, we need to show that (1.2) indeed has a unique solution, and we will carry out this work starting from the next section. For now, assuming the uniqueness (which is postponed to §6), we can quickly complete the proof of Theorem 1.1.
@@
By by $(p-1)p^{*}=p$ and $(q-1)q^{*}=q$, the equality (5.3) is equivalent to $\ell(p^{*},x^{p-1})=\ell(q^{*},y^{q-1})$. Note that the above procedure can be reversed. This proves the equivalence between the equation-system (1.2) and (1.5).
@@
Let $1<p<q<\infty$. We divide the proof into two cases.
@@
With $f_{\widetilde{\rho_{p,q}}}=1+\widetilde{\rho_{p,q}}\chi+\widetilde{\rho_{p,q}}\overline{\chi}$ being a nontrivial extremizer for the modified extremal problem:
@@
It follows that the function $f_{\widetilde{\rho_{p,q}}}=1+\widetilde{\rho_{p,q}}\chi+\widetilde{\rho_{p,q}}\overline{\chi}$ is also an extremal function for the original hypercontractivity inequality. This completes the whole proof of Theorem 1.1 by assuming the uniqueness of the solution to (1.2) inside $(0,1)^{2}$.
@@ 6. Uniqueness of the solution
This section is devoted to the proof of the uniqueness of solution in the open unit square $(0,1)^{2}$ to the equation-system (1.2). While Figures 5 and 7, obtained via symmetrization and blowup techniques, make the uniqueness of the solution in the unit open square appear so self-evident as to seem almost trivial, a rigorous proof of this fact requires a highly nontrivial argument.
@@
Given $1<p<q<\infty$. The equation-system defined by (1.2) has a unique solution in the unit square $(0,1)^{2}=(0,1)\times(0,1)$.
@@
Proof.
By taking derivatives, we obtain
$\frac{\partial h_{1}}{\partial x}$ $=\frac{2(x^{p-1}-1)}{(1+2x)^{2}(1+2x^{p})}\left(\frac{1+2x^{p}}{3}\right)^{\frac{1}{p}}<0,$
$\frac{\partial h_{2}}{\partial x}$ $=\frac{(x^{p-2}-1)(1+2x^{p})+px^{p-2}(x-1)(1+2x)}{(1+2x^{p})^{2}}$
$=\frac{(x^{p-1}-1)(2x^{p-1}+1)+(p-1)x^{p-2}(x-1)(1+2x)}{(1+2x^{p})^{2}}<0.$
This implies that both the abscissa and ordinate of the curve $h(p,\cdot)$ monotonically decrease with respect to $y$. Therefore, the curve $h(p,\cdot)$ has no self-intersecting point, and similarly for $h(q,\cdot)$. It follows that the equivalence in the lemma holds. ∎
@@ 6.1. Symmetrization
Based on the observation $h_{2}(p,x)=h_{2}(p^{*},x^{p-1})$ with $p^{*}=p/(p-1)$, we define a transformation
@@
Having examined symmetry in the domain of $h$, we now turn to symmetry in its range. Let $(u,v)=h(p,x)$. If $p=2$, then $v=3/2-1/(6u^{2})$. Letting $p\to\infty$, we get v=3/21/(2u)v=3/2-1/(2u); letting $p\to 1$, we get $u=1/3$. That is, $p=1,2,\infty$ correspond to $u=1/3$, $u=1/\sqrt{9-6v}$, u=1/(32v)u=1/(3-2v) respectively. See the upper-right corner of Figure 9. Note that $u=1/\sqrt{9-6v}$ is exactly the geometric mean of $u=1/3$ and u=1/(32v)u=1/(3-2v). Define another transformation
@@
We thus observe that $H$ satisfies the following symmetry with respect to $\alpha=0$. Denote $H=(H_{1},H_{2})$. Then
@@ 6.2. Injectivity
$H_{2}\Big(\alpha,\frac{1}{2}\Big)=\frac{(2-2^{\alpha})(2-2^{-\alpha})}{6}\in\Big(0,\ \frac{1}{6}\Big].$
@@
[…]
That is, $\Omega$ is the domain bounded by the dashed line in the bottom-right corner of Figure 9, minus the line segment $L$.
@@
(6.10) $I(0,t)=\frac{1}{t}>\frac{4\log t}{\log\frac{1+2t^{2}}{3}}-2,\quad\forall\ t\in\left(0,\ \frac{1}{2}\right).$
@@
$\frac{\partial I}{\partial\alpha}=\$$\frac{2}{(2-(1+\alpha)t^{1-\alpha}-(1-\alpha)t^{1+\alpha})^{2}}$
$\bigg[\left(4\alpha-(1+\alpha)(t^{1-\alpha}+\frac{1}{t^{1-\alpha}})+(1-\alpha)(t^{1+\alpha}+\frac{1}{t^{1+\alpha}})\right)\log\frac{1}{t}$
$-(\frac{1}{t^{\alpha}}-t^{\alpha})(1-t^{1-\alpha})(1-t^{1+\alpha})\frac{1}{t}\bigg]$
$=:\$$\frac{2(J_{1}\log\frac{1}{t}-J_{2})}{(2-(1+\alpha)t^{1-\alpha}-(1-\alpha)t^{1+\alpha})^{2}}.$
@@
$\frac{\partial J_{1}}{\partial t}=(1-\alpha^{2})(t^{\alpha}-\frac{1}{t^{\alpha}})(1+\frac{1}{t^{2}})<0,$
@@
$\frac{\partial J}{\partial t}J_{1}^{2}=\$$-\frac{\partial J_{2}}{\partial t}J_{1}+\frac{\partial J_{1}}{\partial t}J_{2}-\frac{1}{t}J_{1}^{2}=-\frac{1}{t^{2}}(1-t^{1-\alpha})(1-t^{1+\alpha})$
$\bigg[(t+\frac{1}{t})(t^{\alpha}-\frac{1}{t^{\alpha}})^{2}-2(t-\frac{1}{t})(t^{2\alpha}-\frac{1}{t^{2\alpha}})\alpha$
$+\left((t+\frac{1}{t})(t^{\alpha}+\frac{1}{t^{\alpha}})^{2}+4(t+\frac{1}{t})-8(t^{\alpha}+\frac{1}{t^{\alpha}})\right)\alpha^{2}\bigg].$
@@
$(t+\frac{1}{t})(t^{\alpha}+\frac{1}{t^{\alpha}})^{2}+4(t+\frac{1}{t})-8(t^{\alpha}+\frac{1}{t^{\alpha}})$
$>4(t+\frac{1}{t})-4(t^{\alpha}+\frac{1}{t^{\alpha}})=\frac{4(1-t^{1-\alpha})(1-t^{1+\alpha})}{t}>0.$
@@
$(t-\frac{1}{t})^{2}(t^{\alpha}+\frac{1}{t^{\alpha}})^{2}<(t+\frac{1}{t})\left((t+\frac{1}{t})(t^{\alpha}+\frac{1}{t^{\alpha}})^{2}+4(t+\frac{1}{t})-8(t^{\alpha}+\frac{1}{t^{\alpha}})\right).$
@@
$2(t+\frac{1}{t})(t^{\alpha}+\frac{1}{t^{\alpha}})<(t+\frac{1}{t})^{2}+(t^{\alpha}+\frac{1}{t^{\alpha}})^{2}.$
@@
(6.11) $s(t):=\log\frac{1+2t^{2}}{3}-\frac{4t}{1+2t}\log t<0,\quad\forall\ t\in\left(0,\ \frac{1}{2}\right).$
@@
$r^{\prime}(t)=-\frac{1}{t}\frac{(1-t)(1+2t)(1-2t^{2})}{(1+2t^{2})^{2}}<0,\quad\forall\ t\in\left(0,\ \frac{1}{2}\right).$
@@ 6.3. Separation
$H_{2}(\alpha,t)<H_{2}\left(\alpha,\frac{1}{2}\right)=\frac{5-2(2^{\alpha}+2^{-\alpha})}{6}.$
@@
$I$ $:=24+2\left(2t+\frac{1}{2t}\right)\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)-\left(2t+\frac{1}{2t}-t^{\alpha}-\frac{1}{t^{\alpha}}\right)^{2}.$
@@
These are univariate inequalities involving logarithms, which can be proved directly by methods similar to (6.11) or using the techniques in §B.
@@ 6.4. Transversality
This section focuses on the tangents of the curves in Figure 7. For simplicity, we denote
@@
$=\frac{4}{(1+2t^{2})^{3}}\bigg[(1-\alpha^{2})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\log t+\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)$
$\hskip 60.00009pt+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}}\bigg],$
@@
$=\frac{3t}{4(1+2t^{2})^{2}}\bigg[A\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+C\bigg],$
@@
$I$ $=\frac{4}{(1+2t^{2})^{3}}\left((1-\alpha^{2})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\log t-(1+2t^{2})v_{t}\right),$
@@
$g$ $=\frac{4(1+2t^{2})^{2}}{3t}(J-P)=A\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+C+D(1-\alpha^{2})+E(1-\alpha^{4}),$
@@
$\frac{\partial g}{\partial\alpha}$ $=(A\log t+B)\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\log t-2D\alpha-4E\alpha^{3},$
$\frac{\partial^{2}g}{\partial\alpha^{2}}$ $=(A(\log t)^{2}+2B\log t)\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{2}-2D-12E\alpha^{2},$
$\frac{\partial^{3}g}{\partial\alpha^{3}}$ $=(A(\log t)^{3}+3B(\log t)^{2})\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{3}-24E\alpha,$
$\frac{\partial^{4}g}{\partial\alpha^{4}}$ $=(A(\log t)^{4}+4B(\log t)^{3})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{4}-24E,$
$\frac{\partial^{5}g}{\partial\alpha^{5}}$ $=(A(\log t)^{5}+5B(\log t)^{4})\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{5}.$
@@
(6.21) $\frac{\partial g}{\partial\alpha}(1,t)$ $=(A\log t+B)\left(t-\frac{1}{t}\right)+B\left(t+\frac{1}{t}\right)\log t-2D-4E<0.$
@@
$G$ $=(1-\alpha^{2})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\log t+\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)$
$\ \ \ \ +\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}},$
$\frac{\partial G}{\partial\alpha}$ $=(1-\alpha^{2})\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{2}-\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\log t$
$\ \ \ \ +t^{\alpha}-\frac{1}{t^{\alpha}}+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)\log t,$
$\frac{\partial^{2}G}{\partial\alpha^{2}}$ $=(1-\alpha^{2})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{3}-3\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{2}$
$\ \ \ \ +\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{2},$
$\frac{\partial^{3}G}{\partial\alpha^{3}}$ $=(1-\alpha^{2})\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{4}-5\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{3}$
$\ \ \ \ -3\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{2}+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{3}$
$=\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{3}\left((1-\alpha^{2})\log t+5\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}-\frac{3}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right).$
@@
$\frac{\partial^{3}G}{\partial\alpha^{3}}$ $>\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{3}\left(\log t-\frac{8}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)$
@@
We now verify that $\frac{\partial^{3}G_{k}}{\partial\alpha^{3}}>0$. From the estimate for $\frac{\partial^{3}G}{\partial\alpha^{3}}$ and the fact that $W_{1}^{2}>W_{0}^{2}$, it suffices to prove

### v2

@@ 4.1. Existence of a nontrivial critical extremizer
The existence of a nontrivial critical extremizer is given in the following
@@
By the definition of $r_{p,q}$, we have $G(r_{p,q},\rho)\geq 0$ for all $\rho\in[0,1]$. In particular, $G(r_{p,q},1)\geq 0$. We now prove that $G(r_{p,q},1)\neq 0$. Suppose by contradiction that $G(r_{p,q},1)=0$. Then the left derivative of $G(r_{p,q},\cdot)$ at $\rho=1$ is non-positive. Therefore,
@@
By Lemma 3.1, we may fix a $r_{0}\in(r_{p,q},\sqrt{(p-1)/(q-1)})$. Then by $r_{0}<\sqrt{(p-1)/(q-1)}$ and the expansion (3.2) of $G(r,\rho)$ at $\rho=0$, there exists $\rho_{1}=\rho_{1}(r_{0})\in(0,1)$ such that $G(r_{0},\rho)\geq 0$ for all $\rho\in[0,\rho_{1}]$. Observe that $\|T_{r}f_{\rho}\|_{q}$ is increasing in $r\in[0,1]$, which implies that $G(r,\rho)$ is decreasing with respect to $r$. Hence, for any $r\in[0,r_{0}]$ (and in particular for $r\in(r_{p,q},r_{0}]$), we have
(4.1) $G(r,\rho)\geq G(r_{0},\rho)\geq 0,\ \forall\ \rho\in[0,\rho_{1}].$
Now for any $r\in(r_{p,q},r_{0}]$, by the definition of $r_{p,q}$ and Proposition 3.2, $\min\{G(r,\rho)\mid\rho\in[0,1]\}<0$ and hence
$\min\{G(r,\rho)\mid\rho\in[\rho_{1},1]\}<0,$
since one already has $\min\{G(r,\rho)\mid\rho\in[0,\rho_{1}]\}\geq 0$ by (4.1). Therefore, on the one hand, by the continuity of $G(r,\rho)$, letting $r\to r_{p,q}^{+}$ gives
$\min\{G(r_{p,q},\rho)\mid\rho\in[\rho_{1},1]\}\leq 0,$
and on the other hand, again by the definition of $r_{p,q}$,
$\min\{G(r_{p,q},\rho)\mid\rho\in[\rho_{1},1]\}\geq\min\{G(r_{p,q},\rho)\mid\rho\in[0,1]\}\geq 0.$
It follows that $\min\{G(r_{p,q},\rho)\mid\rho\in[\rho_{1},1]\}=0$. That is, there exists $\rho_{0}\in[\rho_{1},1]$ such that $G(r_{p,q},\rho_{0})=0$. Finally, Lemma 4.2 implies that $\rho_{0}\in[\rho_{1},1)\subset(0,1)$, which finishes the proof. ∎
@@ 4.2. Derivation of the extremal system
To complete the proof of Theorem 1.2, we need to show that (1.4) indeed has a unique solution, and we will carry out this work starting in §6. For now, assuming the uniqueness, we can quickly complete the proof of Theorem 1.2.
@@
By $(p-1)p^{*}=p$ and $(q-1)q^{*}=q$, the equality (5.3) is equivalent to $\ell(p^{*},x^{p-1})=\ell(q^{*},y^{q-1})$. Reversing the above procedure, we obtain the equivalence between (1.4) and (1.7).
@@
Let $1<p<q<\infty$. We give the proof in two cases.
@@
With $f_{\widetilde{\rho_{p,q}}}=1+\widetilde{\rho_{p,q}}\chi+\widetilde{\rho_{p,q}}\overline{\chi}$ being a nontrivial critical extremizer for the modified extremal problem:
@@
It follows that $f_{\widetilde{\rho_{p,q}}}=1+\widetilde{\rho_{p,q}}\chi+\widetilde{\rho_{p,q}}\overline{\chi}$ is also an extremal function for the original hypercontractivity inequality. This completes the whole proof of Theorem 1.2 by assuming the uniqueness of the solution to (1.4) inside $(0,1)^{2}$.
@@ 6. Uniqueness of the solution
This section is devoted to the proof of the uniqueness of solution in the open unit square to (1.4). While Figures 7 and 9, obtained via symmetrization and blowup techniques, make the uniqueness of the solution in $(0,1)^{2}$ appear so self-evident as to seem almost trivial, a rigorous proof of this fact requires a highly nontrivial argument.
@@
Given $1<p<q<\infty$. The equation-system (1.4) has a unique solution in $(0,1)^{2}$.
@@
By taking derivatives, we obtain
$\frac{\partial h_{1}}{\partial x}=\frac{2(x^{p-1}-1)}{(1+2x)^{2}(1+2x^{p})}\left(\frac{1+2x^{p}}{3}\right)^{\frac{1}{p}}<0.$
Since 1x1-x, $1-x^{p-1}$, $1/(1+2x^{p})$ are all strictly decreasing in $x$, we have $\frac{\partial h_{2}}{\partial x}<0$. Therefore both the abscissa and ordinate of the curve $h(p,\cdot)$ monotonically decrease with respect to $x$. In particular, the curve $h(p,\cdot)$ has no self-intersecting point. This implies the following lemma.
@@ 6.1. Symmetrization
Observe that $h_{2}(p,x)=h_{2}(p^{*},x^{p-1})$ with $p^{*}=p/(p-1)$, we define a transformation
@@
Having examined symmetry in the domain of $h$, we now turn to symmetry in its range. Let $(u,v)=h(p,x)$. If $p=2$, then $v=3/2-1/(6u^{2})$. Letting $p\to\infty$, we get v=3/21/(2u)v=3/2-1/(2u); letting $p\to 1$, we get $u=1/3$. That is, $p=1,2,\infty$ correspond to $u=1/3$, $u=1/\sqrt{9-6v}$, u=1/(32v)u=1/(3-2v) respectively. See the upper-right corner of Figure 11. Note that $1/\sqrt{9-6v}$ is exactly the geometric mean of $1/3$ and 1/(32v)1/(3-2v). Define another transformation
@@
Observe that $H$ satisfies the following symmetry with respect to $\alpha=0$. Denote $H=(H_{1},H_{2})$. Then
@@ 6.2. Injectivity
$H_{2}(\alpha,\tfrac{1}{2})=\frac{(2-2^{\alpha})(2-2^{-\alpha})}{6}\in(0,\tfrac{1}{6}].$
@@
$\Omega=\{(u,v)\in\mathbb{R}^{2}\mid 0<v<1,\ \tfrac{1}{2}\log\tfrac{3-2v}{3}<u<-\tfrac{1}{2}\log\tfrac{3-2v}{3}\}\setminus L.$
That is, $\Omega$ is the domain bounded by the dashed line in the bottom-right corner of Figure 11, minus the vertical line segment $L$.
@@
(6.10) $I(0,t)=\frac{1}{t}>\frac{4\log t}{\log\frac{1+2t^{2}}{3}}-2,\quad\forall\ t\in(0,\tfrac{1}{2}).$
@@
$\frac{(2-(1+\alpha)t^{1-\alpha}-(1-\alpha)t^{1+\alpha})^{2}}{2}\ \frac{\partial I}{\partial\alpha}$
$=\Big[4\alpha-(1+\alpha)(t^{1-\alpha}+\tfrac{1}{t^{1-\alpha}})+(1-\alpha)(t^{1+\alpha}+\tfrac{1}{t^{1+\alpha}})\Big]\log\tfrac{1}{t}-(\tfrac{1}{t^{\alpha}}-t^{\alpha})(1-t^{1-\alpha})(1-t^{1+\alpha})\tfrac{1}{t}$
$=:J_{1}\log\tfrac{1}{t}-J_{2}.$
@@
$\frac{\partial J_{1}}{\partial t}=(1-\alpha^{2})(t^{\alpha}-\tfrac{1}{t^{\alpha}})(1+\tfrac{1}{t^{2}})<0,$
@@
$\frac{\partial J}{\partial t}J_{1}^{2}=-\frac{\partial J_{2}}{\partial t}J_{1}+\frac{\partial J_{1}}{\partial t}J_{2}-\frac{1}{t}J_{1}^{2}=-\frac{1}{t^{2}}(1-t^{1-\alpha})(1-t^{1+\alpha})$
$\Big[(t+\tfrac{1}{t})(t^{\alpha}-\tfrac{1}{t^{\alpha}})^{2}-2(t-\tfrac{1}{t})(t^{2\alpha}-\tfrac{1}{t^{2\alpha}})\alpha+\left((t+\tfrac{1}{t})(t^{\alpha}+\tfrac{1}{t^{\alpha}})^{2}+4(t+\tfrac{1}{t})-8(t^{\alpha}+\tfrac{1}{t^{\alpha}})\right)\alpha^{2}\Big].$
@@
$(t+\tfrac{1}{t})(t^{\alpha}+\tfrac{1}{t^{\alpha}})^{2}+4(t+\tfrac{1}{t})-8(t^{\alpha}+\tfrac{1}{t^{\alpha}})>4(t+\tfrac{1}{t})-4(t^{\alpha}+\tfrac{1}{t^{\alpha}})=\frac{4(1-t^{1-\alpha})(1-t^{1+\alpha})}{t}>0.$
@@
$(t-\tfrac{1}{t})^{2}(t^{\alpha}+\tfrac{1}{t^{\alpha}})^{2}<(t+\tfrac{1}{t})\Big[(t+\tfrac{1}{t})(t^{\alpha}+\tfrac{1}{t^{\alpha}})^{2}+4(t+\tfrac{1}{t})-8(t^{\alpha}+\tfrac{1}{t^{\alpha}})\Big].$
@@
$2(t+\tfrac{1}{t})(t^{\alpha}+\tfrac{1}{t^{\alpha}})<(t+\tfrac{1}{t})^{2}+(t^{\alpha}+\tfrac{1}{t^{\alpha}})^{2}.$
@@
(6.11) $s(t):=\log\frac{1+2t^{2}}{3}-\frac{4t}{1+2t}\log t<0,\quad\forall\ t\in(0,\tfrac{1}{2}).$
@@
$r^{\prime}(t)=-\frac{1}{t}\frac{(1-t)(1+2t)(1-2t^{2})}{(1+2t^{2})^{2}}<0,\quad\forall\ t\in(0,\tfrac{1}{2}).$
@@ 6.3. Separation
$H_{2}(\alpha,t)<H_{2}(\alpha,\tfrac{1}{2})=\frac{5-2(2^{\alpha}+2^{-\alpha})}{6}.$
@@
$I$ $:=24+2(2t+\tfrac{1}{2t})(t^{\alpha}+\tfrac{1}{t^{\alpha}})-(2t+\tfrac{1}{2t}-t^{\alpha}-\tfrac{1}{t^{\alpha}})^{2}.$
@@
These are univariate inequalities involving logarithms, which can be proved directly by methods similar to (6.11) or using the techniques in §B in the appendix.
@@ 6.4. Transversality
We turn to the tangents of the curves in Figure 9. For simplicity, denote
@@
Recall from (1.18) that
$B(u,v)=\left(\frac{u}{v^{3/2}},\ v^{1/2}\right).$
@@
$=\frac{4}{(1+2t^{2})^{3}}\Big[(1-\alpha^{2})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\log t+\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}}\Big],$
@@
$=\frac{3t}{4(1+2t^{2})^{2}}\Big[A\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+C\Big],$
@@
$I$ $=\frac{4}{(1+2t^{2})^{3}}\Big[(1-\alpha^{2})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\log t-(1+2t^{2})v_{t}\Big],$
@@
$g$ $=\frac{4(1+2t^{2})^{2}}{3t}(J-P)=A\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+C+D(1-\alpha^{2})+E(1-\alpha^{4}),$
@@
$\frac{\partial g}{\partial\alpha}$ $=(A\log t+B)\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\log t-2D\alpha-4E\alpha^{3},$
$\frac{\partial^{2}g}{\partial\alpha^{2}}$ $=(A(\log t)^{2}+2B\log t)\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{2}-2D-12E\alpha^{2},$
$\frac{\partial^{3}g}{\partial\alpha^{3}}$ $=(A(\log t)^{3}+3B(\log t)^{2})\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}-24E\alpha,$
$\frac{\partial^{4}g}{\partial\alpha^{4}}$ $=(A(\log t)^{4}+4B(\log t)^{3})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{4}-24E,$
$\frac{\partial^{5}g}{\partial\alpha^{5}}$ $=(A(\log t)^{5}+5B(\log t)^{4})\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{5}.$
@@
(6.21) $\frac{\partial g}{\partial\alpha}(1,t)$ $=(A\log t+B)\left(t-\tfrac{1}{t}\right)+B\left(t+\tfrac{1}{t}\right)\log t-2D-4E<0.$
@@
$G$ $=(1-\alpha^{2})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\log t+\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}},$
$\frac{\partial G}{\partial\alpha}$ $=(1-\alpha^{2})\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{2}-\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\log t+t^{\alpha}-\tfrac{1}{t^{\alpha}}+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)\log t,$
$\frac{\partial^{2}G}{\partial\alpha^{2}}$ $=(1-\alpha^{2})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}-3\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{2}+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{2},$
$\frac{\partial^{3}G}{\partial\alpha^{3}}$ $=(1-\alpha^{2})\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{4}-5\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}$
$\ \ \ \ -3\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{2}+\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}$
$=\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}\left((1-\alpha^{2})\log t+5\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}-\frac{3}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right).$
@@
$\frac{\partial^{3}G}{\partial\alpha^{3}}$ $>\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{3}\left(\log t-\frac{8}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)$
@@
We now verify $\frac{\partial^{3}G_{k}}{\partial\alpha^{3}}>0$. By the estimate for $\frac{\partial^{3}G}{\partial\alpha^{3}}$ and the fact that $W_{1}^{2}>W_{0}^{2}$, it suffices to prove